OpenStudy (anonymous):

Please can someone help me. I have to get this done. I'm sick and its kinda hard to focus when ur trying to feel better. my teacher even said this is one of the hardest modules.

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

Subtract and simplify: The quantity 5 x plus 6, all over the binomial x squared minus 1, end quantity, minus the quantity 1 over the binomial x squared minus 1, end quantity.

OpenStudy (anonymous):

gabgurl, where did you get that? I can't see the problem.

OpenStudy (anonymous):

flvs. that is the problem but in word form

OpenStudy (anonymous):

I'm in flvs right now actually........

OpenStudy (anonymous):

\[\frac{ 5x+6 }{ x ^{2}-1}-\frac{ 1 }{ x ^{2}-1 }\]

OpenStudy (anonymous):

really. what subject and whos ur teacher

OpenStudy (anonymous):

OH, I'm sorry I didn't see that you weren't two different people.....

OpenStudy (anonymous):

ok hold on ill try

OpenStudy (anonymous):

since the denominator is the same then:\[\frac{ 5(x+1) }{ (x-1)(x+1) }=\frac{ 5 }{ x-1 }\]

OpenStudy (anonymous):

the first one u wrote ahmad is the one

OpenStudy (anonymous):

wow!

OpenStudy (anonymous):

did I do anything wrong @gabgurl ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so now you have the answer ;)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

later gabgurl, nice to know I'm not the only FLVS girl here! :)

OpenStudy (anonymous):

LOL! byeee

OpenStudy (anonymous):

ahmad do u mind helping me with a couple more? @Ahmad1

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Add and simplify: the quantity 4 over the binomial x minus 7, end quantity, plus the quantity 1 over the binomial x minus 3, end quantity

OpenStudy (anonymous):

multiply by (x-7) and (x-3) to find the LCD you get :\[\frac{ 4(x-3)-(x-7) }{ (x-7)(x-3) }\]

OpenStudy (anonymous):

im still a little confused

OpenStudy (anonymous):

sorry im a pain in the butt

OpenStudy (anonymous):

\[\frac{ 4x-12+x-7 }{ (x-3)(x-7) }=\frac{ 5x-19 }{ (x-3)(x-7) }\] I hope that's right

OpenStudy (anonymous):

no no it's my bad @gabgurl I put - instead of +

OpenStudy (anonymous):

yes that is! thank u

OpenStudy (anonymous):

thank u(: @Ahmad1

OpenStudy (anonymous):

you are so welcome @gabgurl :)

OpenStudy (anonymous):

:)